# template.bib

### v1

@@ 3.3 Perturbation
Perturbation techniques in XAI are used to analyze model sensitivity by altering input data and observing prediction changes [xiongExplainableArtificialIntelligence2024]. We use the perturbation technique from [tempelExplainingHumanActivity2024] to test the two explanation methods and perturb specific parts of the model architecture related to skeleton body key points.
We use GCNs as the backbone of our architectures, which adapt convolutional neural networks to graph data, making them suitable for skeleton-based HAR. In a GCN, nodes aggregate information from neighboring nodes, capturing spatial relationships relevant to human movements. This involves multiplying the node feature matrix by a normalized adjacency matrix $\mathbf{A}$, identity matrix $\mathbf{I}$, and a learnable edge importance matrix $\mathbf{E}$. $\mathbf{A}$ is then split into $\mathbf{A}_{j}$, such that $\mathbf{A}+\mathbf{I}=\sum_{j}\mathbf{A}_{j}$. The output is computed as follows [yanSpatialTemporalGraph2018]:
@@
As in [tempelExplainingHumanActivity2024], we apply the perturbation by modifying edges $e_{n}$ in $\mathbf{E}$ corresponding to the body key points, using the sorted SHAP and GradCAM values as the importance measure to choose which edge has to be perturbed.
@@ 4 Experiments
This section thoroughly describes the experiments conducted on the NTU RGB+D and CP datasets. First, we explain the two datasets and outline how they are preprocessed for the model. Then, we give implementation details regarding the GCN model, GradCAM, and SHAP within our framework. Next, we compare how the explainability techniques map importance onto body key points in the skeleton sequences and the overall feature ranking of those. Further, we perform the perturbation experiments to investigate which of the most important body key points, as determined by SHAP and GradCAM, respectively, lead to a more significant decrease in model performance. Also, we correlate GradCAM values from different layers. Finally, we compare the computation time of both explainability methods within our experimental setting.
@@ 4.2 Implementation
We utilize the best-performing model architecture from [tempelAutoGCNgenericHumanActivity2024] for the NTU RGB+D 60 X-View dataset and the architecture from [tempelLightweightNeuralArchitecture2024] for the CP dataset. In order to compute SHAP values, we apply the ”DeepExplainer“ from the SHAP library [lundbergUnifiedApproachInterpreting2017]. The ”DeepExplainer“ algorithm requires a reference dataset representing the underlying data distribution. For this, a random reference set of $n=100$ instances is sampled, covering all action classes generated for the NTU RGB+D dataset. Similarly, a reference set consisting of $n=100$ randomly selected window samples from both classes is generated for the CP dataset. Since our GPU memory limits the number of background samples that can be processed at once, subsamples of the background ($n=20$) are used in each experiment, with the resulting SHAP values aggregated afterward. Given the hardware accessibility of typical medical and research institutions, we recommend this as a practical approach. The GradCAM implementation utilizes the PyTorch registerhook functionality at the specified layer to acquire the gradients and feature maps [paszkePyTorchImperativeStyle2019]. All experiments are conducted on a single NVIDIA V100 GPU with 32 GB of memory, using the PyTorch framework (version 2.3.1) [paszkePyTorchImperativeStyle2019] and shap (version 0.46.0) [lundbergUnifiedApproachInterpreting2017]. The global random seed for the experiments is set to $1234$.
@@ 4.3 Qualitative Explanation on Skeleton
In Fig. 2 and 3, the results of the SHAP and the GradCAM values are shown when mapped on the skeleton sequence to visualize their contribution at a specific time frame.
For the NTU RGB+D skeleton, which is taken from class 6 (Pick up), it can be observed that the GradCAM activations for the TCN and attention activation layer differ from each other. While the GradCAM values for the TCN layer have a higher portion of activated joints, those from the attention activation layer keep a smaller amount of activated joints. This can be explained by the functionality of the joint attention layer compressing the feature map obtained from the TCN layer to focus on fewer important body key points. Interestingly, the SHAP and GradCAM importance mappings do not agree with each other. For instance, in window $40$, SHAP attributes higher importance to the upper body, particularly the head and shoulder regions. Meanwhile, GradCAM emphasizes the left limbs, giving them a higher GradCAM value. This contrast highlights the differing perspectives of these two XAI techniques on which body regions are most crucial for the model’s predictions.
(a) GradCAM: Temporal convolutional layer
(b) GradCAM: Attention activation layer
@@
Figure 2: Comparison of the spatial explanations on the body key points with GradCAM for the TCN convolutional layer and attention activation layer, and SHAP for class 6 (pick up) of the NTU RGB+D dataset.
For the CP skeleton analysis, a random infant with CP is selected from the validation set $\mathcal{D}_{val}$, and both GradCAM and SHAP values are mapped onto the respective body key points. Notably, the two methods do not always align in their activation of body key points. This discrepancy is evident in window 40, where each method highlights different body regions. However, there is an agreement between SHAP and GradCAM activations in the CP dataset compared to the NTU RGB+D dataset, e.g., in window 10, where both methods give the knees a higher activation. This is potentially due to the smaller window size, which offers more thorough explanations without excessive averaging.
(a) GradCAM: Temporal convolutional layer
(b) GradCAM: Attention activation layer
@@
Figure 3: Comparison of the spatial explanations on the body key points with GradCAM for the TCN convolutional and attention activation layers, and SHAP for an infant with CP. The size and color indicate the activation in the respective body key point.
@@ 4.4 Perturbation
We apply the same perturbation approach as in [tempelExplainingHumanActivity2024]. Hence, we can directly compare the influence of perturbing important or unimportant body key points on the CP and NTU RGB+D datasets. Furthermore, we apply the negated GradCAM values on the TCN layer from the model for the NTU RGB+D dataset. This way, we can also obtain contributions that negatively influence the prediction of the respective class. For the NTU RGB+D dataset, we chose the same perturbation threshold of $0.35\%$ as in [tempelExplainingHumanActivity2024] to be able to compare the GradCAM results against the reported SHAP values from this study. In Fig. 4, we show the influence of perturbing up to 10 key body points on the prediction performance for the NTU RGB+D dataset for three classes.
We compare the perturbation results for the NTU RGB+D dataset when taking the TCN and last activation layers’ gradients before the classifier against SHAP. It can be seen that both explanation methods perform better than perturbing edges randomly for important and unimportant features/gradients. Moreover, perturbing important body key points identified using GradCAM outperform those determined by SHAP across all three classes. The performance drop is more significant when using the key points identified by GradCAM, while SHAP shows a smaller degradation when negating gradients or focusing on unimportant features. While SHAP initially performs on pair or even better than GradCAM, the performance decline stops at around six perturbed joints for class 6 (Pick up) and four for class 16 (Put on a shoe). For the experiment with negated gradients for GradCAM and unimportant features for SHAP - SHAP performs better in class 6 but slightly worse in the two other classes.
@@
We evaluate the perturbation experiment using the Area Under the Curve (AUC) metric for the CP dataset, shown in Fig. 5. The results reveal that perturbation of the TCN layer, as assessed through GradCAM values, causes a sharp decline in the AUC curve, particularly after perturbing four body key points. In contrast, perturbation guided by SHAP explanations leads to a steady AUC decline, implicating a correct determination of the important body key points. On the other hand, the perturbation experiments with the GradCAM values taken after the attention layer from the main branch produce erratic results. This unpredictability can be attributed to the CP model’s architecture, where the attention layer primarily focuses on frames and simultaneously activates all joints for an important window. Consequently, the perturbation experiments, which depend on correct spatial information of the important body key points, result in uncorrelated AUC oscillation. In contrast, the joint attention layer for the NTU RGB+D model concentrates on the important body key points rather than time frames, resulting in a smooth performance decrease with the rising number of perturbed edges.
Figure 5: Perturbation results on the CP dataset. Both XAI methods perform better than randomly perturbing body key points, indicating their reliability. The results when the GradCAMs are computed after the frame attention layer perform poorly, indicating that this layer is not well suited for obtaining spatial information.
@@ 4.5 Body key points ranking
We compare both XAI techniques’ body key point ranking for the NTU RGB+D dataset, shown in Fig. 6. It can be seen that SHAP and GradCAM disagree on the ordering of the body key points for the NTU RGB+D dataset. This is likely due to their conceptual difference, where GradCAM gives a more spatial-orientated explanation. Another potential reason for this differing sorting is the averaging technique, where the GradCAM and SHAP values are averaged over the whole sequence. At the same time, SHAP attributes a more general feature contribution to the interactions with other features for a specific body key point. While GradCAM values have a more extensive spread among the different body key points, SHAP, especially for Class 11, has a smaller interquartile range from the median.
@@
We apply the same body key point ranking for the GradCAM and SHAP values to one infant with CP and one without CP, shown in Fig. 7. Again, it can be observed that the feature rankings do not agree with SHAP and GradCAM values. Interestingly, the GradCAM values from the frame attention and TCN layers do not agree on their body key point ranking. The same reasons as in Sec. 4.4 can be taken to explain this behavior. As with the NTU RGB+D dataset, the SHAP values have a smaller interquartile range from the median, indicating a higher certainty for the ranking. While SHAP gives the highest importance for the infant with CP to the right and left knee with SHAP, the GradCAM values are also the highest in those two body key points for the attention activation layer. For No CP, the most critical features, based on the median value, are the right index finger and right knee, which also align with the TCN GradCAM values.
@@
4.6 GradCAM in Early Layers
To demonstrate the shortcomings of GradCAM in attributing individual input features, we conduct a correlation experiment where GradCAM values are computed for different GCN layers, as shown in Figure 1. We assess the correlation of GradCAM values in 1) initialization layer $L_{\text{in},i}$, 2) TCN layer $L_{\text{TCN},i}$, and, 3) attention activation layer $L_{\text{Att.},i}$ of input branches $i$ to reference values in the attention activation layer of the main branch. We also compute correlation for the sum of input branches ($L_{\text{in},}$, $L_{\text{TCN},}$, and $L_{\text{Att.},}$ respectively) and include in the analysis the TCN layer of the main branch $m$ ($L_{\text{TCN},m}$).
Similar to [dasGradientWeightedClassActivation2022], layerwise GradCAM values are normalized into the range $[0,1]$ before Spearman’s rank correlation coefficient $\rho$ is computed. Our results in Table 1 show that, for the NTU RGB+D dataset and the CP dataset, GradCAM values at input branches do not correlate with reference values at the attention activation layer of the main branch. The lack of correlation at input branches of the GCN suggests that SHAP is superior to GradCAM for attributing individual input features (joint, velocity, bone, and acceleration) and specific input channels.
Table 1: Correlation of GradCAM values obtained from different layers $L$ for the NTU RGB+D and CP datasets.
@@ 4.7 Runtime
Table 2 lists the computation time for GradCAM and SHAP within our environment. We show the time until the individual algorithm arrives at its final explanation. The GradCAM method is computationally faster than the SHAP method for both datasets due to its lower computational complexity and fewer required calculations. GradCAM only necessitates one forward and one backward pass of the obtained gradients. On the contrary, SHAP requires multiple propagations of the input data on every feature input. Additionally, the SHAP computation has to be repeated five times in our environment due to memory restrictions to obtain the representative reference dataset of $n=100$. In contrast, SHAP computes the individual feature contribution for all four input features, while GradCAM only provides the gradients of the chosen layer.
Table 2: Runtime for SHAP and GradCAM explanation on the CP dataset (both classes) and NTU RGB+D X-View (one class) on $\mathcal{D}_{val}$. The GradCAM runtime is for one layer, and the SHAP runtime is for one iteration with the reference dataset $n=20$.
Dataset GradCAM SHAP
@@ 5 Discussion
Our comparison of SHAP and GradCAM within HAR demonstrates that both methods fundamentally differ in their methodology and the outcome of the final explanation. While they both have unique strengths and weaknesses, their applicability depends heavily on specific needs, requirements, and use cases, which we will elaborate on.
Understanding the model’s decision-making process is crucial in CP diagnosis or treatment planning. SHAP can be used to provide a feature-level explanation, giving clinicians insight into how much each biomechanical feature of each body key point contributes to the model’s final decision. This type of information is highly valuable in clinical settings where practitioners need clear, interpretable reasons behind model outputs. For example, knowing that certain features like acceleration or velocity are more influential in a diagnosis might help to guide the focus on these particular features. GradCAM, on the other hand, provides spatial interpretation, which can be helpful in tasks such as reckoning important body key points in the recording at a specific time frame. It highlights regions of high activation in the model’s chosen target layers, which correspond to body key points in the skeletal structure. This explanation is valuable when clinicians need to visualize which body regions are emphasized, though it lacks the granularity of individual feature importance that SHAP provides.
@@
Another important consideration is the chosen convolutional layer for the GradCAM computation in the model. Our experiments indicate that the last convolutional layer within the network preserved the most information relevant to the final decision. This aligns with prior research recommending focusing on the last convolutional layer for maximal interpretability in Grad-CAM [selvarajuGradCAMVisualExplanations2020]. We also showed that the early-layer activations do not correlate with the activations from the TCN or attention layer, which shows the best results in the perturbation experiment. Furthermore, we obtained the GradCAM values from the frame attention layer for the CP dataset, which focuses on the important time frames rather than the body key points. This may activate body key points, which may no longer provide rich spatial information in some windows. This is due to the attention mechanism, which focuses on the important frames and activates them as a whole instead of the important body key points. When using GradCAM for the activation of the attention layer, the explanation should be interpreted in relation to a specific type of attention (e.g., temporal frame attention versus spatial joint attention), and the architecture of the GCN itself should be modified to emphasize the most useful explanation for the end-user.
@@
Within our study, we applied the perturbation approach from [tempelExplainingHumanActivity2024], which may be biased towards a spatial-orientated explanation method: GradCAM. Since GradCAM takes the gradients at the specific body key points without needing a reference dataset like SHAP does, it is superior in this specific experimental setting. GradCAM values reflect activation intensities from the chosen layers, which may not be directly comparable to the SHAP feature-based importance values. For example, Grad-CAM highlights regions based on body key point activations, whereas SHAP reflects how much each feature contributes to the model’s output across the entire reference dataset. Additionally, the perturbation method is applied to the whole video frame for the NTU RGB+D data, with the importance scores averaged over the time frame. While this approach provides a general overview of the particular action’s most crucial body key points, it may obscure critical activations of body key points occurring within shorter temporal windows. By averaging the whole video, we risk overlooking nuanced movements or short-duration actions, which are essential for understanding complex human activities. Future work should investigate how a finer temporal averaging technique and localized perturbation strategies can address this limitation. For instance, perturbing the model over smaller, temporally-defined video segments could provide more granular insights. We might capture more precise and contextually relevant perturbations by classifying shorter video segments and perturbing the network for each segment individually. Such a method could yield more accurate explanations of the model’s behavior, particularly for brief actions involving subtle and short body movements.
@@
SHAP considers feature interactions and how the importance of a body part can change depending on other body parts, while Grad-CAM does not explicitly consider interactions. SHAP provides feature-level attribution, which means it gives importance to the overall contribution of input features and body parts. If a body key point is highly correlated with a particular feature in the model, SHAP might attribute it high importance, regardless of the spatial distribution. SHAP measures how much the body key point contributes to the model’s decision, while Grad-CAM measures how much it contributes to the decision based on its activation. Even if both methods show a body key part as important, the contribution might come from a different context: SHAP values can be influenced by feature interaction and a correlation between the kinematic input features in the model and the underlying reference dataset. In contrast, GradCAM will only consider the direct relationship between the skeleton’s spatial features and the decision. GradCAM hereby examines the regions that activate the most in response to the model’s prediction. It focuses on the body key points where the model’s decision is most influenced, so the resulting heatmap will highlight specific areas that activate the most. However, this might not directly correlate with the feature-level contributions and body key point activations that SHAP captures.
@@
Runtime and computation needs are critical when considering these methods for clinical use. GradCAM’s faster runtime makes it suitable for real-time analysis or applications where speed is essential. SHAP, while slower, provides richer information, making it better suited for offline analysis or in-depth clinical studies where speed and computing resources are less of an issue.
@@
Given these methods’ different strengths and weaknesses, some recommendations can be made for their application in healthcare environments. Grad-CAM might be preferred when a quick, spatial understanding of which body parts the model focuses on is sufficient. This could be useful for real-time diagnostic tools or when working with visual data, such as motion capture or video-based assessments. However, Grad-CAM may miss the ”bigger picture, “mainly when interactions between body key points and their features are crucial to the diagnosis and explanation.

### v2

@@ 3.3 Perturbation
Perturbation techniques in XAI are used to analyze model sensitivity by altering input data and observing prediction changes [xiongExplainableArtificialIntelligence2024]. We use the perturbation technique from [tempelExplainingHumanActivity2025] to test the two explanation methods and perturb specific parts of the model architecture related to skeleton body key points.
We use GCNs as the backbone of our architectures, which adapt convolutional neural networks to graph data [yanSpatialTemporalGraph2018]. With the mentioned advancement over using raw video data as the input, discussed in section 1, GCNs are particularly suitable for skeleton-based HAR. In a GCN, nodes aggregate information from neighboring nodes, capturing spatial relationships relevant to human movements. This involves multiplying the node feature matrix by a normalized adjacency matrix $\mathbf{A}$, identity matrix $\mathbf{I}$, and a learnable edge importance matrix $\mathbf{E}$. $\mathbf{A}$ is then split into $\mathbf{A}_{j}$, such that $\mathbf{A}+\mathbf{I}=\sum_{j}\mathbf{A}_{j}$. The output is computed as follows [yanSpatialTemporalGraph2018]:
@@
As in [tempelExplainingHumanActivity2025], we apply the perturbation by modifying edges $e_{n}$ in $\mathbf{E}$ corresponding to the body key points, using the sorted SHAP and Grad-CAM values as the importance measure to choose which edge has to be perturbed.
@@ 4 Experiments
This section thoroughly describes the experiments conducted on the NTU RGB+D and CP datasets. First, we explain the two datasets and outline how they are preprocessed for the model. Then, we give implementation details regarding the GCN model, Grad-CAM, and SHAP within our framework. Next, we compare how the explainability techniques map importance onto body key points in the skeleton sequences and the overall feature ranking of those. Further, we perform the perturbation experiments to investigate which of the most important body key points, as determined by SHAP and Grad-CAM, respectively, lead to a more significant decrease in model performance. Also, we correlate Grad-CAM values from different layers. Finally, we compare the computation time of both explainability methods within our experimental setting.
@@ 4.2 Implementation
We utilize two GCN models for the comparison, which is the best-performing model architecture from [tempelAutoGCNgenericHumanActivity2024] for the NTU RGB+D 60 X-View dataset and the architecture from [tempelLightweightNeuralArchitecture2024] for the CP dataset. In order to compute SHAP values, we apply the ‘DeepExplainer” from the SHAP library [lundbergUnifiedApproachInterpreting2017]. The “DeepExplainer” algorithm requires a reference dataset representing the underlying data distribution. For this, a random reference set of $n=100$ instances is sampled, covering all action classes generated for the NTU RGB+D dataset. Similarly, a reference set consisting of $n=100$ randomly selected window samples from both classes is generated for the CP dataset. Since our GPU memory limits the number of background samples that can be processed at once, subsamples of the background ($n=20$) are used in each experiment, with the resulting SHAP values aggregated afterward. Given the hardware accessibility of typical medical and research institutions, we recommend this as a practical approach. The Grad-CAM implementation utilizes the PyTorch registerhook functionality at the specified layer to acquire the gradients and feature maps [paszkePyTorchImperativeStyle2019]. All experiments are conducted on a single NVIDIA V100 GPU with 32 GB of memory, using the PyTorch framework (version 2.3.1) [paszkePyTorchImperativeStyle2019] and shap (version 0.46.0) [lundbergUnifiedApproachInterpreting2017]. The global random seed for the experiments is set to $1234$.
@@ 4.3 Qualitative Explanation on Skeleton
In Fig. 2 and 3, the results of the SHAP and the Grad-CAM values are shown when mapped on the skeleton sequence to visualize their contribution at a specific time frame.
For the NTU RGB+D skeleton, which is taken from class 6 (Pick up), it can be observed that the Grad-CAM activations for the TCN and attention activation layer differ from each other. While the Grad-CAM values for the TCN layer have a higher portion of activated joints, those from the attention activation layer keep a smaller amount of activated joints. This can be explained by the functionality of the joint attention layer compressing the feature map obtained from the TCN layer to focus on fewer important body key points. Interestingly, the SHAP and Grad-CAM importance mappings do not agree with each other. For instance, in window $40$, SHAP attributes higher importance to the upper body, particularly the head and shoulder regions. Meanwhile, Grad-CAM emphasizes the left limbs, giving them a higher Grad-CAM value. This contrast highlights the differing perspectives of these two XAI techniques on which body regions are most crucial for the model’s predictions.
(a) Grad-CAM: Temporal convolutional layer
(b) Grad-CAM: Attention activation layer
@@
Figure 2: Comparison of the spatial explanations on the body key points with Grad-CAM for the TCN convolutional layer and attention activation layer, and SHAP for class 6 (pick up) of the NTU RGB+D dataset.
For the CP skeleton analysis, a random infant with CP is selected from the validation set $\mathcal{D}_{val}$, and both Grad-CAM and SHAP values are mapped onto the respective body key points. Notably, the two methods do not always align in their activation of body key points. This discrepancy is evident in window 40, where each method highlights different body regions. However, there is an agreement between SHAP and Grad-CAM activations in the CP dataset compared to the NTU RGB+D dataset, e.g., in window 10, where both methods give the knees a higher activation. This is potentially due to the smaller window size, which offers more thorough explanations without excessive averaging.
(a) Grad-CAM: Temporal convolutional layer
(b) Grad-CAM: Attention activation layer
@@
Figure 3: Comparison of the spatial explanations on the body key points with Grad-CAM for the TCN convolutional and attention activation layers, and SHAP for an infant with CP. The size and color indicate the activation in the respective body key point.
@@ 4.4 Perturbation
We apply the same perturbation approach as in [tempelExplainingHumanActivity2025]. Hence, we can directly compare the influence of perturbing important or unimportant body key points on the CP and NTU RGB+D datasets. Furthermore, we apply the negated Grad-CAM values on the TCN layer from the model for the NTU RGB+D dataset. This way, we can also obtain contributions that negatively influence the prediction of the respective class. For the NTU RGB+D dataset, we chose the same perturbation threshold of $0.35\%$ as in [tempelExplainingHumanActivity2025] to be able to compare the Grad-CAM results against the reported SHAP values from this study. In Fig. 4, we show the influence of perturbing up to 10 key body points on the prediction performance for the NTU RGB+D dataset for three classes.
We compare the perturbation results for the NTU RGB+D dataset when taking the TCN and last activation layers’ gradients before the classifier against SHAP. It can be seen that both explanation methods perform better than perturbing edges randomly for important and unimportant features/gradients. Moreover, perturbing important body key points identified using Grad-CAM outperform those determined by SHAP across all three classes. The performance drop is more significant when using the key points identified by Grad-CAM, while SHAP shows a minor degradation when negating gradients or focusing on unimportant features. While SHAP initially performs on pair or even better than Grad-CAM, the performance decline stops at around six perturbed joints for class 6 (Pick up) and four for class 16 (Put on a shoe). For the experiment with negated gradients for Grad-CAM and unimportant features for SHAP - SHAP performs better in class 6 but slightly worse in the two other classes.
@@
We evaluate the perturbation experiment using the Area Under the Curve (AUC) metric for the CP dataset, shown in Fig. 5. The results reveal that perturbation of the TCN layer, as assessed through Grad-CAM values, causes a sharp decline in the AUC curve, particularly after perturbing four body key points. In contrast, perturbation guided by SHAP explanations leads to a steady AUC decline, implicating a correct determination of the important body key points. On the other hand, the perturbation experiments with the Grad-CAM values taken after the attention layer from the main branch produce erratic results. This unpredictability can be attributed to the CP model’s architecture, where the attention layer primarily focuses on frames and simultaneously activates all joints for an important window. Consequently, the perturbation experiments, which depend on correct spatial information of the important body key points, result in uncorrelated AUC oscillation. In contrast, the joint attention layer for the NTU RGB+D model concentrates on the important body key points rather than time frames, resulting in a smooth performance decrease with the rising number of perturbed edges.
Figure 5: Perturbation results on the CP dataset. Both XAI methods perform better than randomly perturbing body key points, indicating their reliability. The results when the Grad-CAM values are computed after the frame attention layer perform poorly, indicating that this layer is not well suited for obtaining spatial information.
@@ 4.5 Body key points ranking
We compare both XAI techniques’ body key point ranking for the NTU RGB+D dataset, shown in Fig. 6. It can be seen that SHAP and Grad-CAM disagree on the ordering of the body key points for the NTU RGB+D dataset. This is likely due to their conceptual difference, where Grad-CAM gives a more spatial-orientated explanation. Another potential reason for this differing sorting is the averaging technique, where the Grad-CAM and SHAP values are averaged over the whole sequence. At the same time, SHAP attributes a more general feature contribution to the interactions with other features for a specific body key point. While Grad-CAM values have a more extensive spread among the different body key points, SHAP, especially for Class 11, has a smaller interquartile range from the median.
@@
We apply the same body key point ranking for the Grad-CAM and SHAP values to one infant with CP and one without CP, shown in Fig. 7. Again, it can be observed that the feature rankings do not agree with SHAP and Grad-CAM values. Interestingly, the Grad-CAM values from the frame attention and TCN layers do not agree on their body key point ranking. The exact reasons, as in Sec. 4.4, can be taken to explain this behavior. As with the NTU RGB+D dataset, the SHAP values have a smaller interquartile range from the median, indicating a higher certainty for the ranking. While SHAP gives the highest importance for the infant with CP to the right and left knee with SHAP, the Grad-CAM values are also the highest in those two body key points for the attention activation layer. For No CP, the most critical features, based on the median value, are the right index finger and right knee, which also align with the TCN Grad-CAM values.
@@
4.6 Grad-CAM in Early Layers
To demonstrate the shortcomings of Grad-CAM in attributing individual input features, we conduct a correlation experiment where Grad-CAM values are computed for different GCN layers, as shown in Figure 1. We assess the correlation of Grad-CAM values in 1) initialization layer $L_{\text{in},i}$, 2) TCN layer $L_{\text{TCN},i}$, and, 3) attention activation layer $L_{\text{Att.},i}$ of input branches $i$ to reference values in the attention activation layer of the main branch. We also compute correlation for the sum of input branches ($L_{\text{in}}$, $L_{\text{TCN}}$, and $L_{\text{Att.}}$ respectively) and include in the analysis the TCN layer of the main branch $m$ ($L_{\text{TCN},m}$).
Similar to [dasGradientWeightedClassActivation2022], layerwise Grad-CAM values are normalized into the range $[0,1]$ before Spearman’s rank correlation coefficient $\rho$ is computed. Our results in Table 1 show that, for the NTU RGB+D dataset and the CP dataset, Grad-CAM values at input branches do not correlate with reference values at the attention activation layer of the main branch. The lack of correlation at input branches of the GCN suggests that SHAP is superior to Grad-CAM for attributing individual input features (joint, velocity, bone, and acceleration) and specific input channels.
Table 1: Correlation of Grad-CAM values obtained from different layers $L$ for the NTU RGB+D and CP datasets.
@@ 4.7 Runtime
Table 2 lists the computation time for Grad-CAM and SHAP within our environment. We show the time until the individual algorithm arrives at its final explanation. The Grad-CAM method is computationally faster than the SHAP method for both datasets due to its lower computational complexity and fewer required calculations. Grad-CAM only necessitates one forward and one backward pass of the obtained gradients. On the contrary, SHAP requires multiple propagations of the input data on every feature input. Additionally, the SHAP computation has to be repeated five times in our environment due to memory restrictions to obtain the representative reference dataset of $n=100$. In contrast, SHAP computes the individual feature contribution for all four input features, while Grad-CAM only provides the gradients of the chosen layer.
Table 2: Runtime for SHAP and Grad-CAM explanation on the CP dataset (both classes) and NTU RGB+D X-View (one class) on $\mathcal{D}_{val}$. The Grad-CAM runtime is for one layer, and the SHAP runtime is for one iteration with the reference dataset $n=20$.
Dataset Grad-CAM SHAP
@@ 5 Discussion
Our comparison of SHAP and Grad-CAM within HAR demonstrates that both methods fundamentally differ in their methodology and the outcome of the final explanation. While they both have unique strengths and weaknesses, their applicability depends heavily on specific needs, requirements, and use cases, which we will elaborate on.
Understanding the model’s decision-making process is crucial in CP diagnosis or treatment planning. SHAP can be used to provide a feature-level explanation, giving clinicians insight into how much each biomechanical feature of each body key point contributes to the model’s final decision. This type of information is highly valuable in clinical settings where practitioners need clear explanations of the model outputs. For example, knowing that certain features like acceleration or velocity are more influential in a diagnosis might help to guide the focus on these particular features. Grad-CAM provides spatial explanations, which can be helpful in tasks such as reckoning important body key points in the recording at a specific time frame. It highlights regions of high activation in the model’s chosen target layers, which correspond to body key points in the skeletal structure. This explanation is valuable when clinicians need to visualize which body regions are emphasized, though it lacks the granularity of individual feature importance that SHAP provides.
@@
Another important consideration is the chosen convolutional layer for the Grad-CAM computation in the model. Our experiments indicate that the last convolutional layer within the network preserved the most information relevant to the final decision. This aligns with prior research recommending focusing on the last convolutional layer for most informative gradient information [selvarajuGradCAMVisualExplanations2020]. We also showed that the early-layer activations do not correlate with the activations from the TCN or attention layer, which shows the best results in the perturbation experiment. Furthermore, we obtained the Grad-CAM values from the frame attention layer for the CP dataset, which focuses on the important time frames rather than the body key points. This may activate body key points, which may no longer provide rich spatial information in some windows. This is due to the attention mechanism, which focuses on the important frames and activates them as a whole instead of the important body key points. When using Grad-CAM for the activation of the attention layer, the explanation should be interpreted in relation to a specific type of attention (e.g., temporal frame attention versus spatial joint attention), and the architecture of the GCN itself should be modified to emphasize the most useful explanation for the end-user.
@@
Within our study, we applied the perturbation approach from [tempelExplainingHumanActivity2025], which may be biased towards the spatial-orientated explanation method: Grad-CAM. Since Grad-CAM takes the gradients at the specific body key points without needing a reference dataset like SHAP does, it is superior in this specific experimental setting. Grad-CAM values reflect activation intensities from the chosen layers, which may not be directly comparable to the SHAP feature-based importance values. For example, Grad-CAM highlights regions based on body key point activations, whereas SHAP reflects how much each feature contributes to the model’s output across the entire reference dataset. Additionally, the perturbation method is applied to the whole video frame for the NTU RGB+D data, with the importance scores averaged over the time frame. While this approach provides a general overview of the particular action’s most crucial body key points, it may obscure critical activations of body key points occurring within shorter temporal windows. By averaging the whole video, we risk overlooking nuanced movements or short-duration actions, which are essential for understanding complex human activities. Future work should investigate how a finer temporal averaging technique and localized perturbation strategies can address this limitation. For instance, perturbing the model over smaller, temporally-defined video segments could provide more granular insights. We might capture more precise and contextually relevant perturbations by classifying shorter video segments and perturbing the network for each segment individually. Such a method could yield more accurate explanations of the model’s behavior, particularly for brief actions involving subtle and short body movements.
@@
SHAP considers feature interactions and how the importance of a body part can change depending on other body parts, while Grad-CAM does not explicitly consider interactions. SHAP provides feature-level attribution, which means it gives importance to the overall contribution of input features and body parts. If a body key point is highly correlated with a particular feature in the model, SHAP might attribute it to high importance, regardless of the spatial distribution. SHAP measures how much the body key point contributes to the model’s decision, while Grad-CAM measures how much it contributes to the decision based on its activation. Even if both methods show a body key part as important, the contribution might come from a different context: SHAP values can be influenced by feature interaction and a correlation between the kinematic input features in the model and the underlying reference dataset. In contrast, Grad-CAM will only consider the direct relationship between the skeleton’s spatial features and the decision. Grad-CAM hereby examines the regions that activate the most in response to the model’s prediction. It focuses on the body key points where the model’s decision is most influenced, so the resulting heatmap will highlight specific areas that activate the most. However, this might not directly correlate with the feature-level contributions and body key point activations that SHAP captures.
The SHAP explanations depend on the reference dataset [haugBaselinesLocalFeature2021, yuanEmpiricalStudyEffect2023]. In our approach, the background dataset is randomly sampled from $\mathcal{D}_{train}$ for the two datasets, with $n=100$. The variance of $\mathbb{E}[f(x)]$ scales with $\frac{1}{\sqrt{n}}$, so a larger reference dataset $\phi_{0}$ will result in a smaller variance but also results in a more extensive runtime of SHAP [lundbergUnifiedApproachInterpreting2017]. Recognizing the already significantly longer runtime compared to Grad-CAM, a greater background dataset is not feasible for our architecture and the used dataset. Yet, future work should incorporate an informed selection of the reference dataset. This can include choosing videos from the CP dataset, which have different severities of CP and are either positive or negative. Finally, the SHAP values combined with the perturbation experiment should be investigated to determine if they lead to the same reliability outcomes and to which extent the explanations match the randomly chosen reference dataset.
@@
Runtime and computation needs are critical when considering these methods for clinical use. Grad-CAM’s faster runtime makes it suitable for real-time analysis or applications where speed is essential. SHAP, while slower, provides richer information, making it better suited for offline analysis or in-depth clinical studies where speed and computing resources are less of an issue.
@@
Given these methods’ different strengths and weaknesses, some recommendations can be made for their application in healthcare environments. Grad-CAM might be preferred when a quick, spatial understanding of which body parts the model focuses on is sufficient. This could be useful for real-time diagnostic tools or when working with visual data, such as motion capture or video-based assessments. However, Grad-CAM may miss the “bigger picture”, mainly when interactions between body key points and their features are crucial to the diagnosis and explanation.
SHAP inherently satisfies a set of axioms: local accuracy, missingness, and consistency, which provide theoretical guarantees for its explanations based on the Shapley value [lundbergUnifiedApproachInterpreting2017, shapley7ValueNPerson1953]. These axioms ensure that SHAP reliably distributes the model’s output among the input features, excludes non-contributing features, and maintains consistency in attribution when feature contributions change. In [molnarInterpretableMachineLearning2022], the author claims that SHAP might be the only explanation method that satisfies all necessary axioms for reliable model explanation in healthcare. Consequently, these foundational properties make SHAP a more robust and theoretically sound explanation method compared to Grad-CAM, which does not include such axioms.